Sunshine Trading Company (STC) has an expected EBIT of $20,000 in perpetuity, and a tax rate of 40%. Interest rate on STC’s $45,000 debt is 8%, and its WACC is 10%. Blackstone Inc (BI) is an all-equity firm, which is identical to STC, except for the capital structure.

(1) What is the value of STC?                                  

(2) What is the value of BI?                                     

(3) What is the rate of return on BI shares?                                    

(4) What is the rate of return on STC shares?
Answer #1

Solution:-

(1) Since the growth rate in EBIT is zero and its consistent till perpetuity, the value of STC based on its earnings is calculated as follows:

Value of STC= [EBIT*(1-tax rate)]/WACC = [20,000*(1-40%)]/10% = $120,000

Value of STC's shares= Total value of STC firm - debt = $120,000-$45,000= $75,000

So, let's calculate the cost of equity for STC. We know that it has a debt of $45,000 and an equity of $75,000 as calculated above which means that total capital is $120,000.

Cost of debt= Interest rate*(1-tax rate)= 8%*(1-40%) = 4.8%

WACC= 10%

Cost of debt*(debt/total capital) + Cost of equity*(equity/total capital)= 10%

4.8%*(45,000/120,000) + cost of equity*(75,000/120,000)= 10%

1.8% + cost of equity*(0.625)= 10%

Cost of equity= 13.12%

(2) BI has same variables other than the fact that it has only equity and no debt. Since, the businesses are identical, the BI would also have cost of equity same as STC. The cost of equity and thus the WACC for BI therefore is 13.12%

Since the growth rate in EBIT is zero and its consistent till perpetuity, the value of BI based on its earnings is calculated as follows:

Value of BI= [EBIT*(1-tax rate)]/WACC = [20,000*(1-40%)]/13.12% = $91,463.4

(3) The rate of return (ROR) on BI shares is as follows (BI has no interest expense):

ROR= Net income/Value of BI shares = EBIT*(1-tax rate)/Value of BI = 20,000*(1-40%)/91,463.4= 13.12%

(4) The rate of return (ROR) on STC shares is as follows:

Interest expense= $45,000*8%= $3,600

Net income for shareholders= (EBIT- interest)*(1-tax rate)= (20,000-3,600)*(1-40%)= $9,840

ROR= Net income/Value of STC shares = 9,840/75,000= 13.12%